MOLECULAR LINE STUDIES WITH THE GREEN BANK TELESCOPE

\underline{GLEN LANGSTON}, National Radio Astronomy Observatory, Green Bank, WV 24944.

The Green Bank Telescope (GBT) has begun regular observations at 18 to 26.5 GHz, and has begun test observations at 40 to 52 GHz. We summarize the capabilities of the 100m diameter GBT for molecular line studies. High sensitivity molecular line maps are presented for the star forming region Sgr B2 region.